OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

6(s-40) = 5(s-40) 6s – 240 = 5s + 200 6s – 5s = 200 + 240 S = 440 mph in still air this is what I got